OpenStudy (anonymous):

The curve y=ax2 + bx + c passes through the points (1,8),(0,5) and (3,20). Find the values of a,b and c and hence the equation of the curve.

OpenStudy (anonymous):

8 = a*1*1 + b *1 + c 5 = a*0*0 + b*0 + c 20 = a*3*3 + b*3 + c

OpenStudy (anonymous):

You need to solve that system of equations to find the values of a,b and c